# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ C.3 Tabulated AMI results on real-world networks
Method Arch. CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $71.9\pm 3.7$ $\textbf{59.5}\pm 8.2$ $4.8\pm 4.6$ $4.7\pm 8.4$ $0.5\pm 2.5$
MLP $\textbf{74.8}\pm 2.4$ $\underline{58.2}\pm 6.0$ $25.0\pm 3.9$ $41.0\pm 3.0$ $30.7\pm 1.5$
GCN $68.4\pm 2.4$ $54.0\pm 6.4$ $26.5\pm 3.2$ $36.9\pm 4.8$ $40.1\pm 1.7$
GIN $65.2\pm 3.4$ $45.0\pm 6.8$ $24.7\pm 2.6$ $31.5\pm 3.1$ $22.8\pm 3.2$
SAGE $69.0\pm 2.8$ $49.6\pm 6.0$ $27.4\pm 3.6$ $40.4\pm 4.2$ $40.3\pm 1.3$
DMoN LIN $37.6\pm 8.4$ $26.7\pm 14.5$ $1.1\pm 1.4$ $6.2\pm 5.3$ $0.8\pm 1.3$
MLP $55.8\pm 3.4$ $36.7\pm 9.4$ $9.0\pm 2.1$ $24.4\pm 3.0$ $21.5\pm 1.1$
GCN $66.4\pm 1.9$ $48.4\pm 5.7$ $26.6\pm 3.7$ $32.4\pm 3.0$ $36.6\pm 1.0$
GIN $46.9\pm 5.0$ $28.1\pm 7.1$ $16.3\pm 2.7$ $15.8\pm 3.1$ $21.2\pm 3.0$
SAGE $64.7\pm 2.1$ $47.9\pm 5.8$ $26.6\pm 3.7$ $32.6\pm 2.8$ $34.7\pm 0.8$
NOCD LIN $24.5\pm 7.7$ $7.2\pm 4.5$ $3.6\pm 1.2$ $9.4\pm 2.5$ $4.0\pm 1.4$
MLP $\underline{74.0}\pm 1.4$ $56.8\pm 3.9$ $\textbf{39.3}\pm 5.3$ $38.1\pm 2.5$ $28.9\pm 0.5$
GCN $68.0\pm 2.3$ $46.5\pm 7.3$ $24.6\pm 11.2$ $\textbf{42.3}\pm 2.2$ $\textbf{43.6}\pm 0.5$
GIN $60.0\pm 3.8$ $42.4\pm 4.9$ $34.5\pm 3.9$ $27.9\pm 3.2$ $23.7\pm 1.2$
SAGE $69.7\pm 1.9$ $54.5\pm 4.7$ $\underline{36.7}\pm 2.8$ $\underline{41.7}\pm 2.3$ $\underline{42.9}\pm 0.4$
DiffPool LIN $20.0\pm 14.2$ $17.2\pm 18.2$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 OOM
MLP $50.6\pm 6.7$ $30.4\pm 7.9$ $5.1\pm 1.0$ $19.3\pm 3.4$ OOM
GCN $69.0\pm 3.9$ $50.6\pm 6.7$ $22.5\pm 3.1$ $29.3\pm 4.8$ OOM
GIN $28.4\pm 9.1$ $28.2\pm 12.2$ $12.5\pm 3.8$ $5.1\pm 3.0$ OOM
SAGE $65.6\pm 4.5$ $53.6\pm 9.0$ $17.1\pm 2.9$ $33.5\pm 3.2$ OOM
MinCut LIN $68.9\pm 3.0$ $49.1\pm 3.4$ $33.7\pm 3.4$ $19.4\pm 10.4$ $3.8\pm 2.0$
MLP $66.5\pm 1.2$ $44.3\pm 4.8$ $33.2\pm 4.7$ $37.4\pm 2.3$ $21.6\pm 8.1$
GCN $62.1\pm 1.4$ $46.7\pm 5.3$ $28.2\pm 3.2$ $37.5\pm 1.9$ $37.7\pm 1.4$
GIN $60.7\pm 1.4$ $46.1\pm 4.3$ $22.0\pm 2.4$ $24.8\pm 11.9$ $14.2\pm 4.6$
SAGE $62.3\pm 1.2$ $46.2\pm 6.8$ $29.4\pm 3.7$ $37.1\pm 1.9$ $37.3\pm 0.5$
Ortho LIN $57.4\pm 2.6$ $42.6\pm 5.1$ $12.7\pm 3.7$ $18.7\pm 5.0$ $17.1\pm 1.1$
MLP $55.8\pm 2.7$ $29.7\pm 7.4$ $4.3\pm 1.1$ $25.1\pm 1.8$ $17.9\pm 0.5$
GCN $62.2\pm 2.0$ $42.0\pm 5.9$ $19.1\pm 3.2$ $35.7\pm 1.9$ $34.6\pm 0.8$
GIN $51.8\pm 2.7$ $31.7\pm 5.4$ $17.6\pm 2.6$ $23.2\pm 2.6$ $25.5\pm 1.2$
SAGE $60.9\pm 2.1$ $40.1\pm 6.3$ $15.9\pm 3.2$ $34.2\pm 2.2$ $29.5\pm 0.6$
Infomap $40.6\pm 0.2$ $26.7\pm 0.1$ $35.7\pm 0.3$ $37.9\pm 0.2$ $35.9\pm 0.1$
C.4 Tabulated number of detected communities on real-world networks

### v4

@@ D.3 Tabulated AMI results on real-world networks
Method Arch. Cora CiteSeer PubMed PC Photo CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $19.9\pm 10.3$ $12.8\pm 6.9$ $0.9\pm 3.7$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $71.9\pm 3.7$ $\mathbf{59.5\pm 8.2}$ $5.3\pm 9.1$ $11.7\pm 11.0$ $0.0\pm 0$ 0.0
MLP $32.7\pm 4.6$ $13.6\pm 2.6$ $\mathbf{22.9\pm 4.6}$ $33.6\pm 13.3$ $51.3\pm 11.1$ $\mathbf{74.8\pm 2.4}$ $\mathbf{58.2\pm 6.0}$ $24.6\pm 2.9$ $38.7\pm 4.8$ $30.3\pm 1.8$
GCN $31.7\pm 3.0$ $12.8\pm 2.9$ $19.2\pm 5.8$ $\mathbf{49.9\pm 3.6}$ $61.4\pm 4.2$ $68.4\pm 2.4$ $54.0\pm 6.4$ $26.5\pm 2.8$ $36.9\pm 4.2$ $40.2\pm 1.6$
GIN $\mathbf{35.8\pm 4.1}$ $16.1\pm 3.2$ $16.1\pm 5.8$ $40.5\pm 3.9$ $50.3\pm 7.0$ $65.2\pm 3.4$ $45.0\pm 6.8$ $25.0\pm 3.5$ $29.7\pm 8.1$ $22.2\pm 2.5$
SAGE $32.5\pm 5.0$ $12.9\pm 2.7$ $19.8\pm 4.6$ $45.5\pm 5.4$ $49.2\pm 6.1$ $69.0\pm 2.8$ $49.6\pm 6.0$ $26.2\pm 2.8$ $\mathbf{41.0\pm 2.5}$ $39.9\pm 1.6$
DMoN LIN $15.3\pm 11.7$ $9.2\pm 6.9$ $9.1\pm 4.5$ $1.0\pm 0.8$ $2.0\pm 2.1$ $55.1\pm 12.9$ $31.9\pm 19.4$ $18.7\pm 12.0$ $9.1\pm 12.9$ $0.8\pm 1.1$
MLP $\mathbf{36.1\pm 5.2}$ $20.2\pm 3.6$ $19.0\pm 6.5$ $45.0\pm 2.2$ $56.7\pm 3.1$ $72.1\pm 1.1$ $51.5\pm 3.9$ $34.2\pm 3.5$ $38.6\pm 3.0$ $29.7\pm 0.4$
GCN $\mathbf{37.2\pm 4.2}$ $18.4\pm 2.6$ $19.3\pm 4.0$ $46.4\pm 1.9$ $57.7\pm 2.8$ $66.6\pm 1.9$ $48.2\pm 6.1$ $33.8\pm 3.5$ $38.2\pm 2.2$ $40.6\pm 0.4$
GIN $27.4\pm 3.3$ $11.5\pm 3.1$ $10.3\pm 7.1$ $35.3\pm 2.3$ $42.6\pm 4.9$ $53.1\pm 3.7$ $39.5\pm 6.4$ $27.4\pm 2.4$ $22.2\pm 3.5$ $27.1\pm 1.6$
SAGE $\mathbf{37.0\pm 4.0}$ $18.0\pm 3.4$ $17.5\pm 3.7$ $46.4\pm 1.6$ $58.0\pm 3.2$ $66.3\pm 1.6$ $47.2\pm 4.8$ $\underline{35.6\pm 3.8}$ $38.6\pm 2.1$ $40.2\pm 0.4$
NOCD LIN $2.9\pm 1.3$ $3.0\pm 1.6$ $0.2\pm 0.5$ $1.9\pm 2.5$ $2.1\pm 3.0$ $24.5\pm 7.7$ $7.2\pm 4.5$ $3.6\pm 1.2$ $9.4\pm 2.5$ $4.0\pm 1.4$
MLP $\mathbf{33.1\pm 14.7}$ $\mathbf{29.6\pm 10.8}$ $16.8\pm 8.1$ $47.9\pm 2.1$ $55.1\pm 15.4$ $\mathbf{74.0\pm 1.4}$ $\mathbf{56.8\pm 3.9}$ $\mathbf{39.3\pm 5.3}$ $38.1\pm 2.5$ $28.9\pm 0.5$
GCN $32.6\pm 4.6$ $11.7\pm 1.7$ $\mathbf{19.3\pm 7.3}$ $\mathbf{49.6\pm 2.2}$ $\mathbf{64.4\pm 4.1}$ $68.0\pm 2.3$ $46.5\pm 7.3$ $24.6\pm 11.2$ $\mathbf{42.3\pm 2.2}$ $\mathbf{43.6\pm 0.5}$
GIN $\mathbf{35.1\pm 4.8}$ $19.2\pm 3.4$ $17.0\pm 6.3$ $32.6\pm 3.7$ $43.3\pm 4.1$ $60.0\pm 3.8$ $42.4\pm 4.9$ $34.5\pm 3.9$ $27.9\pm 3.2$ $23.7\pm 1.2$
SAGE $32.5\pm 3.7$ $11.6\pm 2.8$ $\mathbf{22.6\pm 4.9}$ $47.7\pm 2.4$ $\mathbf{63.3\pm 3.5}$ $69.7\pm 1.9$ $54.5\pm 4.7$ $\underline{36.7\pm 2.8}$ $\mathbf{41.7\pm 2.3}$ $\underline{42.9\pm 0.4}$
DiffPool LIN $0.8\pm 1.9$ $1.2\pm 0.5$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $20.0\pm 14.2$ $17.2\pm 18.2$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 OOM
MLP $4.9\pm 1.3$ $4.1\pm 1.4$ $10.3\pm 3.6$ $1.6\pm 0.4$ $3.5\pm 1.4$ $50.6\pm 6.7$ $30.4\pm 7.9$ $5.1\pm 1.0$ $19.3\pm 3.4$ OOM
GCN $28.5\pm 4.1$ $12.4\pm 2.5$ $18.8\pm 5.5$ $30.4\pm 9.1$ $46.4\pm 6.5$ $69.0\pm 3.9$ $50.6\pm 6.7$ $22.5\pm 3.1$ $29.3\pm 4.8$ OOM
GIN $14.0\pm 4.8$ $8.2\pm 3.1$ $2.1\pm 1.7$ $5.3\pm 5.6$ $5.7\pm 6.8$ $28.4\pm 9.1$ $28.2\pm 12.2$ $12.5\pm 3.8$ $5.1\pm 3.0$ OOM
SAGE $24.4\pm 3.0$ $11.6\pm 2.5$ $18.9\pm 5.0$ $3.1\pm 0.6$ $12.8\pm 6.1$ $65.6\pm 4.5$ $53.6\pm 9.0$ $17.1\pm 2.9$ $33.5\pm 3.2$ OOM
MinCut LIN $30.6\pm 4.3$ $16.7\pm 2.6$ $\mathbf{21.0}\pm 1.9$ $3.1\pm 8.5$ $1.9\pm 6.4$ $68.9\pm 3.0$ $49.1\pm 3.4$ $33.7\pm 3.4$ $19.4\pm 10.4$ $3.8\pm 2.0$
MLP $30.4\pm 4.5$ $15.5\pm 3.5$ $\mathbf{19.7\pm 6.3}$ $38.8\pm 13.3$ $56.4\pm 2.7$ $66.5\pm 1.2$ $44.3\pm 4.8$ $33.2\pm 4.7$ $37.4\pm 2.3$ $21.6\pm 8.1$
GCN $31.9\pm 3.7$ $13.1\pm 2.8$ $18.5\pm 4.2$ $46.3\pm 2.0$ $56.1\pm 2.9$ $62.1\pm 1.4$ $46.7\pm 5.3$ $28.2\pm 3.2$ $37.5\pm 1.9$ $37.7\pm 1.4$
GIN $\mathbf{35.7\pm 2.9}$ $18.9\pm 3.0$ $20.2\pm 4.1$ $32.4\pm 7.8$ $49.2\pm 4.9$ $60.7\pm 1.4$ $46.1\pm 4.3$ $22.0\pm 2.4$ $24.8\pm 11.9$ $14.2\pm 4.6$
SAGE $33.6\pm 2.9$ $13.9\pm 3.5$ $17.1\pm 3.7$ $43.7\pm 2.5$ $54.0\pm 2.8$ $62.3\pm 1.2$ $46.2\pm 6.8$ $29.4\pm 3.7$ $37.1\pm 1.9$ $37.3\pm 0.5$
Ortho LIN $8.2\pm 1.2$ $8.0\pm 2.3$ $15.7\pm 4.6$ $12.4\pm 6.5$ $13.6\pm 8.6$ $57.4\pm 2.6$ $42.6\pm 5.1$ $12.7\pm 3.7$ $18.7\pm 5.0$ $17.1\pm 1.1$
MLP $4.4\pm 1.0$ $3.5\pm 1.0$ $8.5\pm 3.6$ $22.4\pm 5.7$ $27.1\pm 2.9$ $55.8\pm 2.7$ $29.7\pm 7.4$ $4.3\pm 1.1$ $25.1\pm 1.8$ $17.9\pm 0.5$
GCN $22.0\pm 3.0$ $10.2\pm 2.1$ $16.4\pm 4.8$ $45.1\pm 1.9$ $54.4\pm 3.6$ $62.2\pm 2.0$ $42.0\pm 5.9$ $19.1\pm 3.2$ $35.7\pm 1.9$ $34.6\pm 0.8$
GIN $19.2\pm 3.6$ $9.4\pm 2.9$ $6.3\pm 3.0$ $37.4\pm 3.8$ $45.9\pm 3.0$ $51.8\pm 2.7$ $31.7\pm 5.4$ $17.6\pm 2.6$ $23.2\pm 2.6$ $25.5\pm 1.2$
SAGE $21.5\pm 4.1$ $7.6\pm 2.5$ $15.9\pm 3.8$ $29.4\pm 2.3$ $36.6\pm 3.2$ $60.9\pm 2.1$ $40.1\pm 6.3$ $15.9\pm 3.2$ $34.2\pm 2.2$ $29.5\pm 0.6$
Infomap $35.2\pm 0.2$ $\underline{23.6\pm 0.1}$ $16.0\pm 0.1$ $\mathbf{49.5\pm 0.5}$ $57.3\pm 1.1$ $40.6\pm 0.2$ $26.7\pm 0.1$ $\underline{35.7\pm 0.3}$ $37.9\pm 0.2$ $35.9\pm 0.1$
D.4 Significance of best Neuromap results vs. best baseline
Table 7: Independent two-sample t-test between 25 samples of AMI values for Neuromap vs. the best baseline for 512 hidden features and $s=\left|Y\right|$. The p-values indicate for which datasets the null hypothesis “the samples have the same mean” can be rejected. Red p-values highlight cases where the best baseline performs significantly better than Neuromap.
Dataset Best Neuromap Best Baseline p
Cora GIN $35.8\pm 4.1$ DMoN GCN $37.2\pm 4.2$ $0.25$
CiteSeer GIN $16.1\pm 3.2$ NOCD MLP $29.6\pm 10.8$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.4e^{-06}}}$
PubMed MLP $22.9\pm 4.6$ NOCD SAGE $22.6\pm 4.9$ $0.83$
PC GCN $49.9\pm 3.6$ NOCD GCN $49.6\pm 2.2$ $0.73$
Photo GCN $61.4\pm 4.2$ NOCD GCN $64.4\pm 4.1$ $0.02$
CS MLP $74.8\pm 2.4$ NOCD MLP $74.0\pm 1.4$ $0.15$
Physics LIN $59.5\pm 8.2$ NOCD MLP $56.8\pm 3.9$ $0.17$
Cora ML GCN $26.5\pm 2.8$ NOCD MLP $39.3\pm 5.3$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.6e^{-12}}}$
Wiki CS SAGE $41.0\pm 2.5$ NOCD GCN $42.3\pm 2.2$ $0.05$
ogb-arxiv GCN $40.2\pm 1.6$ NOCD GCN $43.6\pm 0.5$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.8e^{-11}}}$
D.5 Significance of Neuromap results vs. Infomap
Table 8: Independent two-sample t-test between 25 samples of AMI values for Neuromap with different (G)NN architectures vs. Infomap for $512$ hidden features and $s=\left|Y\right|$. The p-values indicate when the null hypothesis “the samples have the same mean” can be rejected. Blue p-values highlight cases where Neuromap performs significantly better than Infomap; red p-values highlight cases where Infomap performs significantly better than Neuromap.
Dataset LIN MLP GCN GIN SAGE
Cora ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.6e^{-7}}}$ $0.01$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{9.2e^{-6}}}$ $0.47$ $0.02$
CiteSeer ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{6.4e^{-8}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.0e^{-15}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.1e^{-15}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.7e^{-11}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.1e^{-16}}}$
PubMed ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.2e^{-16}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.5e^{-7}}}$ $0.01$ $0.95$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{4.9e^{-4}}}$
PC ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.8e^{-48}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.1e^{-6}}}$ $0.55$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.4e^{-11}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.6e^{-3}}}$
Photo ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.8e^{-42}}}$ $0.02$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{6.3e^{-5}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.7e^{-5}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.1e^{-6}}}$
CS ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{6.2e^{-24}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.8e^{-29}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{3.0e^{-27}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.5e^{-22}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{8.8e^{-26}}}$
Physics ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{3.1e^{-16}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{6.5e^{-19}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.3e^{-17}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.6e^{-12}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.2e^{-16}}}$
Cora ML ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.7e^{-14}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.3e^{-16}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.5e^{-14}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{9.9e^{-14}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{6.5e^{-15}}}$
Wiki CS ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.4e^{-11}}}$ $0.42$ $0.29$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.3e^{-5}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{3.4e^{-6}}}$
ogb-arxiv ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{3.7e^{-62}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.5e^{-14}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.0e^{-12}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.3e^{-19}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{5.4e^{-12}}}$
D.6 Tabulated number of detected communities on real-world networks